OpenStudy (anonymous):

Hulk left home and walked 8 blocks west. Then he turned and walked 6 blocks north. If each block is 500 ft. long, how far is Hulk from home?.. Please Show Work ;]

OpenStudy (anonymous):

you're finding hypotenuse (8*500)^2+(6*500)^2=c^2 16000000+9000000=25000000 distance is 5000